Suppose an investment offers to quintuple your money in 36 months (don’t believe it). What rate of return per quarter are you being offered? (Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/400)^4n
where
A=future value($5x)
P=present value($x say)
r=rate of interest
n=time period.(36 months=3 years)

5x=x*(1+r/400)^(4*3)

(5)^(1/12)=(1+r/400)

(1+r/400)=1.143529836

r=1.143529836-1

=14.35%(Approx).=rate of return per quarter.